OpenStudy (anonymous):

simplify 3-2[3(x+2)-7]+x.

OpenStudy (cornzyblack):

3 - [2*[3(x+2) - 7]] + x = 3 - [6(x+2) - 14] + x = 3 - [(6x + 12) - 14] + x = 3 - [6x +12 - 14] + x= 3 - [6x -2} + x = 3 - 6x +2 + x = collecting like terms, 3+2 - 6x + x= 5 -5x = 5(1-x)